OpenStudy (anonymous):

trigonometric identities? I'm given the identity sin(x) x cot(x) = 1, and am told to solve for cos(x), but how do I find the number value? any help would be great. thanks in advance! edit: i'm having the same problem with trying to figure out this one i'm given cos(x) = cot(x), and told to find csc.

OpenStudy (shaik0124):

see we know that tanx =1/cotx now cotx=1/tanx=cosx/sinx from above and we know that tanx=sinx/cosx so now replace this in sinx x cosx=1 we will get sinx x cosx/sinx=1 sinx gets cancelled off so cosx=1

OpenStudy (anonymous):

@shaik0124 thank you! for the second one, i understand that cos = b/c, and cot = b/a. csc = c/a, but how do i plug in numbers to solve the identity?

OpenStudy (shaik0124):

for 2nd question u cna solve like this cosx=cosx/sinx cosx=cosx x 1/sinx cosx=cosx x cscx where cscx=1/sinx divide by cosx on both sidees 1=cscx

OpenStudy (anonymous):

thank you so much.

OpenStudy (shaik0124):

understood@mee